Consider a particle of mass m confined in an infinite 1-D potential well of width a;
V (x)=0 for 0 less or equal to x and x less than or equal to a,
V(x)=infinite otherwise
Find the eigenstates of the Hamiltonian and the corresponding eigen energies Using three possibilities considering the energy E,
E>0
E<0
E=0
1
Expert's answer
2019-09-10T13:25:43-0400

The eigenvalues of the Hamiltonian are:


"H|\psi\rangle=E|\psi\rangle"

for every eigenvector.

Write the Hamiltonian:


"H=-\frac{\hbar^2}{2m}\frac{\text{d}^2}{\text{d}x^2},"

substitute this to the first equation:


"-\frac{\hbar^2}{2m}\frac{\text{d}^2}{\text{d}x^2}|\psi\rangle=E|\psi\rangle,"

the solution of this equation gives the set of eigenstates:


"|\psi\rangle=C\space\text{sin}\Big(\frac{\pi n x}{a}\Big)."

The corresponding energies (eigenvalues):


"E=\frac{n^2\pi^2\hbar^2}{2ma}."